# Typical thermalization of low-entanglement states

### v1

@@ Abstract
$D_{l}(g_{\beta}(H),g_{\beta}(UHU^{\dagger}))\leq C_{3}\,N^{-\gamma_{4}(1-\alpha)}$ (15)
for system-size independent constants $C_{3},\gamma_{4}$. The proof may be found in the Supplemental Material, Section II, and easily generalizes to other choices of $\delta$. As anticipated, the unitary ensemble $\mathcal{E}(\delta)$ is chosen in order to approximately preserve the energy of any state; this implies that $U\sim\mathcal{E}(\delta)$ approximately commutes with the Hamiltonian, and we show $\|H-UHU^{\dagger}\|_{\infty}\leq\delta$. Taking $\delta$ as in Theorem 2, this leads to the following result
$\|e^{-iHt}\rho e^{iHt}-e^{-iUHU^{\dagger}t}\rho e^{iUHU^{\dagger}t}\|_{1}\leq 2tN^{\frac{1-\alpha}{D+1}-\kappa},$ (16)
which means that the dynamics under $U$ and $UHU^{\dagger}$ are indistinguishable up to a time $t^{*}\sim N^{\kappa-\frac{1-\alpha}{D+1}}$. If $\kappa>0$, that is the BE error decays faster than the worst-case scenario, $\alpha$ can be chosen such that $t^{*}$ increases with the system size. In other words $H$ and $UHU^{\dagger}$ generate nearly the same dynamics for a time $t^{*}\sim\mathrm{poly}(N)$.
It is now worth noting that if both $\rho$ and $\sigma$ are translation invariant, the averaging over regions in the definition of $D_{l}(\rho,\sigma)$ can be dropped, making the indistinguishability statement valid for any observable supported on an individual small region. We show that if the original Hamiltonian is translation-invariant, we recover this property to some extent in the equilibrium state of the perturbed Hamiltonian. More specifically, we show that if $O$ is an observable supported in $C\in\mathcal{C}_{l}$ and $H$ is translation invariant, for $U$ drawn from $\mathcal{E}(\delta)$, with probability at least $1-\sum_{k}\exp(-c{d_{k}\epsilon}/{\|O\|_{\infty}})$ for a constant $c$ where $d_{k}=\mathrm{dim}(\mathcal{W}_{k})$, we have
$\left|\mathrm{tr}\left(\left(g_{\beta}(H)-\rho_{\infty}^{UHU^{\dagger}}\right)O\right)\right|\leq\epsilon+D_{l}\left(\rho_{\infty}^{UHU^{\dagger}},g_{\beta}(H)\right).$ (17)
@@
Turning to notions of dynamical thermalization, we have investigated the typical time-evolution of the expectation value of a generic observable $A$, $\langle A\rangle_{\rho}:=\mathrm{tr}\left(A\rho\right)$, under the evolution generated by $UHU^{\dagger}$. In the Supplemental Material, Section III, we show that, with probability at least $1-\sum_{k}\exp\left(-c{d_{k}\epsilon}/{\|O\|_{\infty}}\right)$ for a constant $c$, the time- evolution is bounded by
@@
$R(t)\sim\|A\|_{\infty}\frac{N^{2}}{\delta^{4}t^{2}}.$ (19)
@@
In this work, we have made substantial progress in the long-standing research quest of proving thermalization from first principles of quantum mechanics, by showing thermalization of low-entanglement states under typical Hamiltonians. In particular, we have defined ensembles which naturally emerge from the time evolution of low-entanglement states under typical Hamiltonian evolution and show that they are locally indistinguishable from the Gibbs ensemble. The typicality in the Hamiltonian is given by randomizing the eigenbasis locally in energy, and we show that this randomization does not affect the Gibbs state locally. Furthermore, we show that if the Gibbs state has a relatively generic smoothness property, the randomization does not affect the short time dynamics in any discernible way.
On a higher level, this is the one of the main physical messages of this work: it is unrealistic to assume that a Hamiltonian can be specified up to arbitrary precision. The ubiquity of thermalization may then be explained by considering that its absence requires instead a high precision in the specification of the system. In this light, proving thermalization for all, or a large class of, local Hamiltonians might not be required for explaining this observation.
@@
Several natural open questions remain, importantly, related to identifying precise properties of the Hamiltonian achieving a lower than worst case BE error. For these Hamiltonians, our result predicts that a vast ensemble of other Hamiltonians exist that have indistinguishable short time dynamics and which eventually thermalize. Furthermore, while our randomization achieves thermalization, more work may be needed about exploring the precise physical mechanisms that can be held responsible for them. On the one hand, being able to implement such transformations in a controlled way might have implications in the form of algorithms for Gibbs state preparation. On the other, understanding in what condition an uncertainty in either state or Hamiltonian preparation can translate into a random energy-preserving perturbation such as the one we defined might shed light on the stability of many-body localization under realistic conditions. It is the hope that this work contributes to understanding that under a “trembling hand”, most systems follow the laws of quantum statistical mechanics.
@@ Acknowledgements.
We thank Henrik Wilming and Ingo Roth on discussions during early stages of this project. This work has been supported by the DFG (FOR 2724 and CRC 183), the BMBF, the FQXi, and the ERC (DebuQC).

### v2

@@ Abstract
$D_{l}(g_{\beta}(H),g_{\beta}(UHU^{\dagger}))\leq O(N^{-\gamma_{4}\alpha-\gamma_{4}\kappa})$ (15)
for system-size independent constants $\gamma_{4},\gamma_{5}$. The proof may be found in the Supplemental Material, Section II, and easily generalizes to other choices of $\delta$. As anticipated, the unitary ensemble $\mathcal{E}(\delta)$ is chosen in order to approximately preserve the energy of any state; this implies that $U\sim\mathcal{E}(\delta)$ approximately commutes with the Hamiltonian, and we show $\|H-UHU^{\dagger}\|_{\infty}\leq\delta$. For the choice of $\delta$ as in Theorem 2 we derive the following result
$\|\mathrm{e}^{-iHt}\rho\mathrm{e}^{iHt}-\mathrm{e}^{-iH^{\prime}t}\rho\mathrm{e}^{iH^{\prime}t}\|_{1}\leq 2t\,O\left(N^{\frac{1-\alpha}{D+1}-\kappa}\right),$ (16)
with $H^{\prime}=UHU^{\dagger}$. This means that the dynamics under $U$ and $UHU^{\dagger}$ are indistinguishable up to a time $t^{*}\sim N^{\kappa-\frac{1-\alpha}{D+1}}$. If $\kappa>0$, that is the BE error decays faster than the worst-case scenario, $\alpha$ can be chosen such that $t^{*}$ increases with the system size. In other words $H$ and $UHU^{\dagger}$ generate nearly the same dynamics for a time $t^{*}\sim\mathrm{poly}(N)$. Finally, we would like to emphasize that our rigorous approach allows to put on precise and solid ground some of the results obtained on equilibration in Ref. [43].
It is now worth noting that if both $\rho$ and $\sigma$ are translation-invariant, the averaging over regions in the definition of $D_{l}(\rho,\sigma)$ can be dropped, making the indistinguishability statement valid for any observable supported on an individual small region. We show that if the original Hamiltonian is translation-invariant, we recover this property to some extent in the equilibrium state of the perturbed Hamiltonian. More specifically, consider an observable $A$ supported in $C\in\mathcal{C}_{l}$ and $H$ translation-invariant. For $U$ drawn from $\mathcal{E}(\delta)$ we show that, with probability at least $1-\sum_{k}\exp(-c{d_{k}\epsilon}/{\|A\|^{2}_{\infty}})$ for a constant $c$, it holds, with $d_{k}=\mathrm{dim}(\mathcal{W}_{k})$,
$\left|\mathrm{tr}\left(\left(g_{\beta}(H)-\rho_{\infty}^{UHU^{\dagger}}\right)A\right)\right|\leq\epsilon+D_{l}\left(\rho_{\infty}^{UHU^{\dagger}},g_{\beta}(H)\right).$ (17)
@@
Turning to notions of dynamical thermalization, we have investigated the typical time-evolution of the expectation value of a generic observable $A$, $\langle A\rangle_{\rho}:=\mathrm{tr}\left(A\rho\right)$, under the evolution generated by $UHU^{\dagger}$. In the Supplemental Material, Section III, we show that, with probability at least $1-\sum_{k}\exp\left(-c{d_{k}\epsilon}/{\|A\|^{2}_{\infty}}\right)$ for a constant $c$, the time-evolution is bounded by
@@
$R(t)\sim\|A\|_{\infty}\frac{N^{2}}{\delta^{2}t^{2}}.$ (19)
@@
In this work, we have made substantial progress in the long-standing research quest of proving thermalization from first principles of quantum mechanics, by showing thermalization of low-entanglement states under typical Hamiltonians. In particular, we have defined ensembles which naturally emerge from the time-evolution of low-entanglement states under typical Hamiltonian evolution and show that they are locally indistinguishable from the Gibbs ensemble. The typicality in the Hamiltonian is given by randomizing the eigenbasis locally in energy, and we show that this randomization does not affect the Gibbs state locally. Furthermore, we show that if the Gibbs state has a relatively generic smoothness property, the randomization does not affect the short-time dynamics in any discernible way.
On a higher level, this is one of the main physical messages of this work: it is unrealistic to assume that a Hamiltonian can be specified up to arbitrary precision. The ubiquity of thermalization may then be explained by considering that its absence requires instead a high precision in the specification of the system. In this light, proving thermalization for all, or a large class of, local Hamiltonians might not be required for explaining this observation.
@@
Several natural open questions remain. Importantly, the problem of identifying precise properties of the Hamiltonian achieving a lower than worst case BE error. For these Hamiltonians, our result predicts that a vast ensemble of other Hamiltonians exists that all have indistinguishable short time dynamics and which eventually thermalize. Furthermore, while our randomization achieves thermalization, more work may be needed about exploring the precise physical mechanisms that can be held responsible for them. On the one hand, being able to implement such transformations in a controlled way might have implications in the form of algorithms for Gibbs state preparation. On the other hand, understanding under what condition an uncertainty in either state or Hamiltonian preparation can translate into a random energy-preserving perturbation such as the one we defined might shed light on the stability of many-body localization under realistic conditions. It is the hope that this work contributes to understanding that under a “trembling hand”, most systems follow the laws of quantum statistical mechanics.
@@ Acknowledgements.
We thank Henrik Wilming and Ingo Roth on discussions during early stages of this project. This work has been supported by the DFG (FOR 2724 and CRC 183), the BMBF (MUNIQCAtoms), the Quantum Flagship (PasQuans2), the FQXi, and the ERC (DebuQC).